pls help me ,I couldnt solve these


q1: if y=ex+e^x+e^x+...........∞,then dy/dx equals


q2:if x=at2 and y=2at then (xy*d2y/dx2)  is equal to


thank u

Grade:10

2 Answers

AKASH GOYAL AskiitiansExpert-IITD
13 years ago

Dear Rishabh

I think there is some mistake in question 1

confirm the qustion and try to do it by taking log

question2

dy/dx= (dy/dt)/(dx/dt)

dy/dt=2a

dx/dt=2at

dy/dx=1/t

d2y/dx2= d/dx(dy/dx)

use chain rule

d2y/dx2= d/dt(dy/dx)*dt/dx

=d/dt(1/t)*1/2at

=-(1/t2)*1/2at

=-1/2at3

xyd2y/dx2=at2*2at*(-1/2at3)= -a

 

 All the best.                                                          

AKASH GOYAL

AskiitiansExpert-IIT Delhi

 

Please feel free to post as many doubts on our discussion forum as you can. We are all IITians and here to help you in your IIT JEE preparation.

Win exciting gifts by answering the questions on Discussion Forum. So help discuss any query on askiitians forum and become an Elite Expert League askiitian.

Now you score 5+15 POINTS by uploading your Pic and Downloading the Askiitians Toolbar  respectively : Click here to download the toolbar..

 

 

 

 

 

 

 

vikas askiitian expert
13 years ago

1)

y=ex+e^x+e^x+...........∞

Y =ex+y 

dy/dx = ex+y(1+dy/dx)

dy/dx = (ex+y/1-ex+y )


2)

x=at2 ...........1

y=2at    ..............2

x/y = t/2   .........3

put t in eq 2

y = 4ax/y

y2 = 4ax

now differentiating above eq

2ydy/dx = 4a         .............1

again differentiating

2[(dy/dx)2 +yd2y/dx2] = 0        

d2y/dx2 = -(dy/dx)2/y          ..........2

multiply above eq by xy

(d2y/dx2)xy = -x(dy/dx)2

                      =-x(2a/y)2                    (from eq 1)

approve if u like my ans